skip to content
ATNF VLBI Wiki
User Tools
Admin
Log In
Site Tools
Search
Tools
Show pagesource
Old revisions
Export to PDF
Backlinks
Recent Changes
Media Manager
Sitemap
Log In
>
Recent Changes
Media Manager
Sitemap
You are here:
start
»
lbaops
»
lbafeb2008
»
todo
Trace:
•
todo
Sidebar
VLBI Schedule
LBA Home Page
Overview
Documents
Observing
LBA Correlation
Active proposals
eVLBI
Observer's Wiki
Schedules
Sensitivity Calculator
Other VLBI
Edit sidebar
lbaops:lbafeb2008:todo
Parkes
Test Apple Xserve RAIDS
DONE
Connect DAS to pkvsi1 (not v1pks)
Check VSIC connections
Test DAS
Mopra
Copy DAS profiles from v1mop
DONE
Reconfig network setup
DONE
Get turret rotation/translation working again
Verify 12mm Rx setup/operation
Verify all updated software working.
Setup hybrid for circular polarisation at 12mm. May have to be done on Wedn.
ATCA
'
Reconnect VSIC and test
DONE
Test dodgey channel on VSIC/cables
DONE
Hobart
Check CVS version
DONE
Ceduna
Check CVS version
Check can connect to green
General
Create VC fringes test experiments
Check CVS versio upto date
Merge UDP version?
Check cdisko working OK
Investigate cdisko scheduling
Check RtFC works with green
Copy schedules to all observatories and run atcasched at Mopra/ATCA
Label all disks
Update the disk status page on the wiki with disk IDs
lbaops/lbafeb2008/todo.txt
· Last modified: 2015/12/18 16:38 by
127.0.0.1
Page Tools
Show pagesource
Old revisions
Backlinks
Export to PDF
Back to top